# Improving Matrix Exponential for Generative AI Flows: A Taylor-Based Approach Beyond Paterson–Stockmeyer

### v1

@@ 3.3 Proposed algorithms
17:  $p=2\cdot i-1$
18:  if $m=1$ then
19:   $E_{1}=C_{p}\cdot\|{W}\|_{1}^{2}$
20:   $E_{2}=C_{p+1}\cdot\|{W}\|_{1}^{3}$
21:  else
22:   $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}$
23:   $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}$
24:   if $j\cdot k=m$ then
25:     $E_{1}=E_{1}\cdot\|{W}\|_{1}$
26:     $E_{2}=E_{2}\cdot\|{W}^{2}\|_{1}$
27:   else
28:     $E_{2}=E_{2}\cdot\|{W}\|_{1}$
@@
30:  end if
31:  if $E_{1}+E_{2}\leq\varepsilon$ then
32:   $finish=1$
33:  else
34:   $i=i+1$
35:  end if
36:end while
37:if $finish=0$ then
38:  for $i=1$ to 22 do
39:   $s1=\left\lceil{\log_{2}(E_{i}/\varepsilon)/(m+i)}\right\rceil$
40:   if $s1>s$ then
41:     $s=s1$
42:   end if
43:  end for
44:end if
45:if $s>20$ then
46:  $s=20$
47:end if
48:return $m,s$
@@ 4.1 Numerical results in MATLAB
The numerical experiments are conducted against two primary benchmarks: the Taylor-based approach described in [25] (Algorithm 1) and a custom implementation of the Paterson–Stockmeyer (P–S) scheme. This selection is predicated on the current state of deep learning frameworks such as PyTorch, where native matrix exponential routines (e.g., torch.linalg.expm) are typically optimized for fixed IEEE precision formats and do not support user-defined variable error tolerances. Since [25] serves as the established reference for matrix exponential applications within generative flow models, it constitutes the appropriate baseline for assessing improvements in this specific domain. Furthermore, the inclusion of the P–S scheme as a secondary benchmark allows for a direct quantification of the efficiency gains attributable specifically to the advanced polynomial evaluation formulas integrated into the proposed algorithm.
@@
    expm_flow_ps: An implementation based on the proposed Algorithm 2, where matrix polynomials are evaluated using the Paterson–Stockmeyer technique.
@@
    expm_flow_sastre: A version identical to the previous implementation, but utilizing evaluation formulas (10)-(17) for matrix polynomials.
@@
Graphically, Figure 1 shows the results obtained from the first of the experiments, carried out on the matrices comprising MCT and EMP. Figure 1(a) displays the normwise relative error committed by each of the codes for the different matrices analysed. These same errors, ordered from highest to lowest for code expm_flow_sastre, are also shown in Figure 1(b). The black line appearing in 1(a) corresponds to the expected relative error when calculating the exponential of the matrices. This value is equivalent to the product of the condition number of the exponential function for each matrix multiplied by the error tolerance used ($10^{-8}$) in our experiments. Ideally, it is to be expected that all results will be below this line, with those ones further down clearly being better. It is easy to see how the results for expm_flow are less accurate than those for expm_flow_sastre and expm_flow_ps, as they occupy the highest positions in the graph.
@@
Figure 1(c) corresponds to the performance profile. It represents, for each value $\alpha$ on the $x$-axis, the percentage of matrices, expressed as a proportion of one, for which the error incurred in calculating their exponential function is $\alpha$ times less than or equal to the minimum error committed by each of the methods compared. The higher the values provided, the better the code. As we can appreciate, expm_flow_sastre and expm_flow_ps occupy the top of the picture, while expm_flow provides the poorest results. While expm_flow_ps offers better results for $\alpha$ equal to 1, this improvement is reversed in favour of expm_flow_sastre from $\alpha$ values close to 4.
@@
The direct correspondence of these values for $\alpha$ equal to 1 are those ones collected, in the form of a pie chart, on the left side of Figure 1(d). It represents the percentage of cases in which each code yielded the most accurate result. Thus, we can state that expm_flow_ps provided the closest values to the exact solution in 58% of cases, expm_flow_sastre in 42% of matrices, and expm_flow in a negligible percentage close to 0. In contrast, the right side of the picture reports the percentage of matrices for which each method gave the most inaccurate result. As we can see, expm_flow accounted for 99% of the matrices, clearly being the code with the worst performance.
@@
As whisker plots, Figures 1(e) and Figure 1(f) display respectively the order of the polynomial applied and the value of the parameter $s$ used in the scaling and squaring technique. In each box, the red centre line indicates the median, and the blue bottom and top lines indicate the 25th and 75th percentiles, correspondingly. The whiskers extend to the most extreme data points that are not considered outliers, and outliers are represented individually by the “+” marker symbol. The code expm_flow_ps usually worked with polynomials of degree 16, or 15 in the case of expm_flow_sastre, while expm_flow used polynomials of lower order, in most cases between 5 and 7. Regarding the scaling parameter $s$, the median value was 1 for the code expm_flow_ps, 2 for expm_flow_sastre, and 5 for expm_flow. The maximum values reached were equal to 20 for expm_flow_ps and expm_flow_sastre and 718 for expm_flow.
@@
Finally, Figures 1(g) and 1(h) present, in the form of a bar chart, two key comparative results: the number of matrix products required by each code and the response times (expressed in seconds) employed in the computation of all test matrices. Specifically, expm_flow_ps required 3,110 products, which is approximately $1.20$ times the 2,597 products needed by the baseline expm_flow_sastre. The least efficient code, expm_flow, required 5,397 matrix products, representing a factor of approximately $2.08$ compared to expm_flow_sastre. This comparison highlights expm_flow_sastre as the most product-efficient implementation, followed by expm_flow_ps, with expm_flow consistently requiring over twice the computational effort in terms of matrix multiplications.
@@
The analysis of execution times confirms the relative performance difference, using expm_flow_sastre as the baseline for comparison. Specifically, expm_flow_ps required $1.56$ seconds for computation, making it approximately $1.13$ times slower than the $1.39$ seconds required by expm_flow_sastre. The least efficient code, expm_flow, required $2.31$ seconds, resulting in a time factor of approximately $1.67$ times slower compared to expm_flow_sastre. This comparison confirms expm_flow_sastre is the fastest implementation, with expm_flow exhibiting the largest runtime overhead. Notably, these execution time ratios are lower than the corresponding matrix product ratios. This reduced proportionality occurs because the test matrices, which range in size from $4\times 4$ to $1024\times 1024$, are not all large enough for the number of computed matrix products to fully dominate the total execution cost.
@@ 4.2 Performance evaluation in the PyTorch generative flow environment
In this subsection, the results of tests performed directly on the code available in [25] will be described. As mention before, this code incorporates a matrix exponential function into neural networks implementing generative flows models. In our executions, the matrix exponential is computed using the following functions implemented in PyTorch:
@@
    expm_flow_ps is based on Algorithm 2 and evaluates the polynomial using the Paterson–Stockmeyer method.
@@
    expm_flow_sastre: Analogous to the previous function, this implementation is based on Algorithm 2, but the polynomials are evaluated using formulas (10)-(17).
@@
All executions were carried out using CUDA and the graphics card of the machine involved. As testbeds in training, CIFAR-10 [37], ImageNet32 and ImageNet64 [38] datasets were used. The number of invocations analysed to extract results to the functions responsible for calculating the matrix exponential during the training process was 5000. For each call to the exponential function, multiple data were recorded for the matrices comprising the tensor to be computed, such as the number of matrices, their sizes, and the highest matrix $\infty$-norm of all of them. Thus, we can confirm that the $\infty$-norm of the matrices studied varied from 2.84e-4 to 12.57 for CIFAR-10, from 1.17e-05 to 12.49 for ImageNet32, and from 1.27e-05 to 12.8 for ImageNet64.
@@
Figures 2, 3, and 4 graphically illustrate the results obtained for CIFAR-10, ImageNet32 and ImageNet64 datasets, respectively. Figures 2(a), 3(a), and 4(a) plot the normwise relative errors just for the first 150 training runs. For clarity, these same results are shown sorted, but now for the 5000 runs, from largest to smallest error independently for the three functions analysed in Figures 2(b), 3(b), and 4(b). Unlike 1(b), there is here no correspondence between a matrix on the $x$-axis and the three errors that appear on the $y$-axis, due to the aforementioned sorting. As we can observe, expm_flow always resulted in higher errors in the exponential approximation. In contrast, expm_flow_sastre and expm_flow_ps provided very similar results, with expm_flow_sastre being slightly more accurate for the three datasets. This conclusion can also be deduced from the Figures 2(c), 3(c), and 4(c) related to the performance profile, where expm_flow_sastre and expm_flow_ps quickly provided probabilities equal to 1.
Figures2(d), 3(d), and 4(d) also confirm that code expm_flow_sastre delivered a more accurate result than the other functions in a higher percentage of cases, with values of 50%, 49%, and 50% of the matrices for the three datasets, respectively. In turn, function expm_flow_sastre provided similar but slightly lower percentages, with values of 48%, for CIFAR-10, and 45%, for ImageNet32 and ImageNet64.
@@
Finally, with regard to the number of products, Figures 2(g), 3(g), and 4(g) state that expm_flow carried out 1.99, 1.86, and 1.88 times more products than expm_flow_sastre in the calculation of matrix exponentials. Similarly, Figures 2(h), 3(h), and 4(h) show that the exponential computation based on function expm_flow_sastre was 1.87, 1.97, and 2.5 times faster than the same calculation using function expm_flow.
@@ 5 Performance analysis in generative flow models
To evaluate the efficiency gains of the proposed matrix exponential approach (expm_flow_sastre) relative to the baseline (expm_flow), we conducted a series of numerical experiments using three standard computer vision datasets: CIFAR-10, ImageNet32, and ImageNet64. All tests were executed on a system equipped with an Intel Xeon Platinum 8268 Processor, an NVIDIA RTX8000 GPU and 128 GB of RAM.
The evaluation consists of two parts. First, isolated benchmarks were performed to compare the execution times of both methods across various matrix sizes, independent of other neural network components. Second, both algorithms were integrated into the coupling layers of a Glow-based architecture [39], following the methodology described in [25]. Performance is quantified by measuring total computation time and calculating the speed-up ratio, defined as the execution time of expm_flow divided by that of expm_flow_sastre.

### v2

@@ 3.3 Proposed algorithms
17:  if $m=2$ then $W^{2}=W^{j-1}\cdot W$
18:  end if
19:  $p=2\cdot i-1$
20:  if $m=1$ then
21:    $E_{1}=C_{p}\cdot\|{W}\|_{1}^{2}$
22:    $E_{2}=C_{p+1}\cdot\|{W}\|_{1}^{3}$
23:  else if $j\cdot k=m$ then % m=2, 4, 8
24:    $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}\|_{1}$
25:    $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}^{2}\|_{1}$
26:  else % m=15
27:    $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}$
28:    $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}\|_{1}$
@@
30:  if $E_{1}+E_{2}\leq\varepsilon$ then
31:    $finish=1$
32:  else
33:    $i=i+1$
34:  end if
35:end while
36:if $finish=0$ then
37:  for $i=1$ to 22 do
38:    $s1=\left\lceil{\log_{2}(E_{i}/\varepsilon)/(m+i)}\right\rceil$
39:    if $s1>s$ then
40:     $s=s1$
41:    end if
42:  end for
43:end if
44:if $s>20$ then
45:  $s=20$
46:end if
47:return $m,s$
$M$ Computed $W$ Power Required Norms Estimated Norms
1 $W$ $\|W^{2}\|_{1}$ $\|W^{3}\|_{1}$ $\|W\|^{2}_{1}$ $\|W\|^{3}_{1}$
2 $W^{2}$ $\|W^{3}\|_{1}$ $\|W^{4}\|_{1}$ $\|W^{2}\|^{1}_{1}\cdot\|W\|_{1}$ $\|W^{2}\|^{1}_{1}\cdot\|W^{2}\|_{1}$
4 $W^{2}$ $\|W^{5}\|_{1}$ $\|W^{6}\|_{1}$ $\|W^{2}\|^{2}_{1}\cdot\|W\|_{1}$ $\|W^{2}\|^{2}_{1}\cdot\|W^{2}\|_{1}$
8 $W^{2}$ $\|W^{9}\|_{1}$ $\|W^{10}\|_{1}$ $\|W^{2}\|^{4}_{1}\cdot\|W\|_{1}$ $\|W^{2}\|^{4}_{1}\cdot\|W^{2}\|_{1}$
even $m$ $W^{j}$ $\|W^{m+1}\|_{1}$ $\|W^{m+2}\|_{1}$ $\|W^{j}\|^{k}_{1}\cdot\|W\|_{1}$ $\|W^{j}\|^{k}_{1}\cdot\|W^{2}\|_{1}$
15 $W^{2}$ $\|W^{16}\|_{1}$ $\|W^{17}\|_{1}$ $\|W^{2}\|^{8}_{1}$ $\|W^{2}\|^{8}_{1}\cdot\|W\|_{1}$
odd $m$ $W^{j}$ $\|W^{m+1}\|_{1}$ $\|W^{m+2}\|_{1}$ $\|W^{j}\|^{k}_{1}$ $\|W^{j}\|^{k}_{1}\cdot\|W\|_{1}$
Table 5: Computed powers of $W$, the norms required by (42), and their corresponding approximations for each polynomial degree $M$ considered in Algorithm 3.3. Note that $j=2$ for all degrees $m>1$.
@@ 4.1 Numerical results in MATLAB
The numerical experiments are evaluated against two primary benchmarks: the Taylor-based method described in [23] (Algorithm 1) and a custom implementation of the Paterson–Stockmeyer (P–S) scheme. This selection is predicated on the current state of deep learning frameworks such as PyTorch, where native matrix exponential routines (e.g., torch.linalg.expm) are typically optimized for standard IEEE precision formats and do not support user-defined error tolerances. To evaluate the performance of our proposed methods in a realistic setting, we adopt the framework from [23] as a baseline for matrix exponential applications in generative flow models. Furthermore, the inclusion of the P–S scheme as a secondary benchmark allows for a direct quantification of the efficiency gains attributable specifically to the optimized polynomial evaluation formulas integrated into the proposed algorithm.
@@
    expm_flow_ps: An implementation based on the proposed Algorithms  2, where matrix polynomials are evaluated using the Paterson–Stockmeyer technique, and $3.$
@@
    expm_flow_opt15: A version identical to the previous implementation, but utilizing evaluation formulas (10)-(17) for matrix polynomials. Algorithm 3.3 is now considered.
@@
Graphically, Figure 1 shows the results obtained from the first of the experiments, carried out on the matrices comprising MCT and EMP. Figure 1(a) displays the normwise relative error committed by each of the codes for the different matrices analyzed. These same errors, ordered from highest to lowest for code expm_flow_opt15, are also shown in Figure 1(b). The black line appearing in 1(a) corresponds to the expected relative error when calculating the exponential of the matrices. This value is equivalent to the product of the condition number of the exponential function for each matrix multiplied by the error tolerance used ($10^{-8}$) in our experiments. Ideally, it is to be expected that all results will be below this line, with those ones further down clearly being better. It is easy to see how the results for expm_flow are less accurate than those for expm_flow_opt15 and expm_flow_ps, as they occupy the highest positions in the graph.
@@
Figure 1(c) corresponds to the performance profile. It represents, for each value $\alpha$ on the $x$-axis, the percentage of matrices, expressed as a proportion of one, for which the error incurred in calculating their exponential function is $\alpha$ times less than or equal to the minimum error committed by each of the methods compared. The higher the values provided, the better the code. As we can appreciate, expm_flow_opt15 and expm_flow_ps occupy the top of the picture, while expm_flow provides the poorest results. While expm_flow_ps offers better results for $\alpha$ equal to 1, this improvement is reversed in favour of expm_flow_opt15 from $\alpha$ values close to 4.
@@
The direct correspondence of these values for $\alpha$ equal to 1 are those ones collected, in the form of a pie chart, on the left side of Figure 1(d). It represents the percentage of cases in which each code yielded the most accurate result. Thus, we can state that expm_flow_ps provided the closest values to the exact solution in 58% of cases, expm_flow_opt15 in 42% of matrices, and expm_flow in a negligible percentage close to 0. In contrast, the right side of the picture reports the percentage of matrices for which each method gave the most inaccurate result. As we can see, expm_flow accounted for 99% of the matrices, clearly being the code with the worst performance.
@@
As whisker plots, Figures 1(e) and Figure 1(f) display respectively the order of the polynomial applied and the value of the parameter $s$ used in the scaling and squaring technique. In each box, the red centre line indicates the median, and the blue bottom and top lines indicate the 25th and 75th percentiles, correspondingly. The whiskers extend to the most extreme data points that are not considered outliers, and outliers are represented individually by the “+” marker symbol. The code expm_flow_ps usually worked with polynomials of degree 16, or 15 in the case of expm_flow_opt15, while expm_flow used polynomials of lower order, in most cases between 5 and 7. Regarding the scaling parameter $s$, the median value was 1 for the code expm_flow_ps, 2 for expm_flow_opt15, and 5 for expm_flow. The maximum values reached were equal to 20 for expm_flow_ps and expm_flow_opt15 and 718 for expm_flow.
@@
Finally, Figures 1(g) and 1(h) present, in the form of a bar chart, two key comparative results: the number of matrix products required by each code and the response times (expressed in seconds) employed in the computation of all test matrices. Specifically, expm_flow_ps required 3110 products, which is approximately $1.20$ times the 2597 products needed by the baseline expm_flow_opt15. The least efficient code, expm_flow, required 5397 matrix products, representing a factor of approximately $2.08$ compared to expm_flow_opt15. This comparison highlights expm_flow_opt15 as the most product-efficient implementation, followed by expm_flow_ps, with expm_flow consistently requiring over twice the computational effort in terms of matrix multiplications.
@@
The analysis of execution times confirms the relative performance difference, using expm_flow_opt15 as the baseline for comparison. Specifically, expm_flow_ps required $1.56$ seconds for computation, making it approximately $1.13$ times slower than the $1.39$ seconds required by expm_flow_opt15. The least efficient code, expm_flow, required $2.31$ seconds, resulting in a time factor of approximately $1.67$ times slower compared to expm_flow_opt15. This comparison confirms expm_flow_opt15 is the fastest implementation, with expm_flow exhibiting the largest runtime overhead. Notably, these execution time ratios are lower than the corresponding matrix product ratios. This reduced proportionality occurs because the test matrices, which range in size from $4\times 4$ to $1024\times 1024$, are not all large enough for the number of computed matrix products to fully dominate the total execution cost.
@@ 4.2 Performance evaluation in the PyTorch generative flow environment
In this subsection, the results of tests performed directly on the code available in [23] will be described. As mention before, this code incorporates a matrix exponential function into neural networks implementing generative flow models. In our executions, the matrix exponential is computed using the following functions implemented in PyTorch:
@@
    expm_flow_ps is based on Algorithms 2 and $3.$ The polynomial is evaluated using the Paterson–Stockmeyer method.
@@
    expm_flow_opt15: Analogous to the previous function, this implementation is based on Algorithms 2 and 3.3, where the polynomials are evaluated using formulas (10)-(17).
@@
To maintain consistency with the baseline implementation expm_flow [23], both expm_flow_ps and expm_flow_opt15 utilize the $\infty$-norm.
All computations were performed using NVIDIA CUDA on a dedicated GPU. The proposed method was evaluated during the training process using the CIFAR-10 [35], ImageNet32, and ImageNet64 [36] datasets. To obtain representative performance metrics, we analyzed 5000 invocations of the matrix exponential function. For each invocation, we recorded metadata for the matrices comprising the input tensor, specifically the number of matrices in the batch, their dimensions, and the maximum $\infty$-norm across the set.
The $\infty$-norms of the matrices studied range from $2.84\times 10^{-4}$ to $12.57$ for CIFAR-10, from $1.17\times 10^{-5}$ to $12.49$ for ImageNet32, and from $1.27\times 10^{-5}$ to $12.8$ for ImageNet64. For a normal matrix with the minimum observed norm of $1.27\times 10^{-5}$, the baseline Algorithm 1 requires a Taylor polynomial of degree $m=7$ with a cost of $6M$ matrix products according to (7). In contrast, Algorithm 3.3 utilizes a degree $m=8$ with a cost of only $3M$. Since the scaling parameter is $s=0$ for both, the baseline method theoretically requires twice as many matrix products as the proposed approach.
Similarly, for a normal matrix with the maximum observed norm of $12.8$, the baseline Algorithm 1 selects $m=8$ and $s=5$, leading to a total cost of $12M$ matrix products. In contrast, Algorithm 3.3 utilizes $m=15$ with $s=3$, which reduces the cost to $7M$. This implies that the baseline method is theoretically $1.71$ times more expensive than the proposed approach for this case.
@@
Figures 2, 3, and 4 graphically illustrate the results obtained for CIFAR-10, ImageNet32 and ImageNet64 datasets, respectively. Figures 2(a), 3(a), and 4(a) plot the normwise relative errors just for the first 150 training runs. For clarity, these same results are shown sorted, but now for the 5000 runs, from largest to smallest error independently for the three functions analyzed in Figures 2(b), 3(b), and 4(b). Unlike 1(b), there is here no correspondence between a matrix on the $x$-axis and the three errors that appear on the $y$-axis, due to the aforementioned sorting. As we can observe, expm_flow always resulted in higher errors in the exponential approximation. In contrast, expm_flow_opt15 and expm_flow_ps provided very similar results, with expm_flow_opt15 being slightly more accurate for the three datasets. This conclusion can also be deduced from the Figures 2(c), 3(c) and 4(c) related to the performance profile, where expm_flow_opt15 and expm_flow_ps quickly provided probabilities equal to 1.
Figures2(d), 3(d), and 4(d) also confirm that code expm_flow_opt15 delivered a more accurate result than the other functions in a higher percentage of cases, with values of 50%, 49%, and 50% of the matrices for the three datasets, respectively. In turn, function expm_flow_opt15 provided similar but slightly lower percentages, with values of 48%, for CIFAR-10, and 45%, for ImageNet32 and ImageNet64.
@@
Finally, with regard to the number of products, Figures 2(g), 3(g), and 4(g) state that expm_flow carried out 1.99, 1.86, and 1.88 times more products than expm_flow_opt15 in the calculation of matrix exponentials. Similarly, Figures 2(h), 3(h), and 4(h) show that the exponential computation based on function expm_flow_opt15 was 1.87, 1.97, and 2.5 times faster than the same calculation using function expm_flow.
@@ 5 Performance analysis in generative flow models
To evaluate the efficiency gains of the proposed matrix exponential approach (expm_flow_opt15) relative to the baseline (expm_flow), we conducted a series of numerical experiments on the three aforementioned datasets: CIFAR-10, ImageNet32, and ImageNet64. All tests were executed on a system equipped with an Intel Xeon Platinum 8268 Processor, an NVIDIA RTX8000 GPU and 128 GB of RAM.
Note that in this section both the matrix exponential and the reduced-rank approximation (8) are used following the generative model framework in [23]. Specifically, the matrix exponential is utilized for matrices smaller than $24\times 24$, while the approximation (8) is applied to larger matrices. In the context of the reduced-rank approximation, where no scaling is applied ($s=0$), the computational cost of the baseline algorithm increases linearly with the degree $m$ according to (7). Consequently, the proposed algorithm is significantly more efficient, as demonstrated in Section 3.1 for Taylor based evaluation formula with approximation order $m=15+$. Furthermore, this performance gap widens as the polynomial degree increases. Moreover, as suggested by Remark 1, these efficiency ratios could be more significant for nonnormal matrices.
The evaluation consists of two parts. First, isolated benchmarks were performed to compare the execution times of both methods across various matrix sizes, independent of other neural network components. Second, both algorithms were integrated into the coupling layers of a Glow-based architecture [37], following the methodology described in [23]. Performance is quantified by measuring total computation time and calculating the speed-up ratio, defined as the execution time of expm_flow divided by that of expm_flow_opt15.